### v1

# CuriGS: Curriculum-Guided Gaussian Splatting for Sparse View Synthesis

Abstract
3D Gaussian Splatting (3DGS) has recently emerged as an efficient, high-fidelity representation for real-time scene reconstruction and rendering. However, extending 3DGS to sparse-view settings remains challenging because of supervision scarcity and overfitting caused by limited viewpoint coverage. In this paper, we present CuriGS, a curriculum-guided framework for sparse-view 3D reconstruction using 3DGS. CuriGS addresses the core challenge of sparse-view synthesis by introducing student views: pseudo-views sampled around ground-truth poses (teacher). For each teacher, we generate multiple groups of student views with different perturbation levels. During training, we follow a curriculum schedule that gradually unlocks higher perturbation level, randomly sampling candidate students from the active level to assist training. Each sampled student is regularized via depth-correlation and co-regularization, and evaluated using a multi-signal metric that combines SSIM, LPIPS, and an image-quality measure. For every teacher and perturbation level, we periodically retain the best-performing students and promote those that satisfy a predefined quality threshold to the training set, resulting in a stable augmentation of sparse training views. Experimental results show that CuriGS outperforms state-of-the-art baselines in both rendering fidelity and geometric consistency across various synthetic and real sparse-view scenes. Project page: https://zijian1026.github.io/CuriGS/
Figure 1: CuriGS is a curriculum-guided Gaussian Splatting framework that enhances sparse-view 3D reconstruction by progressively introducing pseudo-views with increasing perturbations, yielding stable geometry and photorealistic rendering from extremely limited input views.
1 Introduction
In recent years, 3D reconstruction has advanced rapidly, becoming a key enabling technology in fields such as virtual reality, digital twins, cultural heritage preservation, and intelligent manufacturing [yu2024sensors, tao2022digital, skublewska20223d]. Among various approaches [schonberger2016structure, wang2018pixel2mesh, qi2017pointnet], Neural Radiance Fields (NeRF) [mildenhall2021nerf] and its variants [barron2021mip, somraj2023vip, chen2021mvsnerf, pumarola2021d] have achieved impressive results in high-fidelity reconstruction and photorealistic view synthesis using implicit volumetric representations, albeit with slow optimization and heavy computational costs. Recently, 3D Gaussian Splatting (3DGS) [kerbl20233d] has emerged as a more efficient and stable explicit representation, enabling real-time rendering by modeling scenes as point-based Gaussians with covariance, color, and opacity attributes [qian20243dgs, liu2024citygaussian, wang2024view, mallick2024taming, zhang2025wheat3dgs, szymanowicz2024splatter]. Beyond representation efficiency, sparse view synthesis [long2022sparseneus, han2024binocular, cai2024structure, xu2024grm, charatan2024pixelsplat] has become a key research focus, aiming to reconstruct accurate geometry and realistic appearance from only a few input images. This setting is both practical and highly challenging, as limited viewpoints lead to supervision scarcity and severe overfitting.
In response to these challenges, recent studies have explored various strategies for enhancing 3D reconstruction under sparse-input conditions. Within the NeRF family, SparseNeRF [wang2023sparsenerf] introduces weak supervision by distilling depth-ordering priors, thereby constraining volumetric geometry and alleviating degradation under extremely limited viewpoints. FreeNeRF [yang2023freenerf] approaches the problem from a frequency and occlusion regularization perspective, imposing simple yet effective constraints on NeRF’s frequency spectrum and near-camera density distribution, which improves stability and generalization for few-shot rendering without external supervision.
In the context of 3DGS, FSGS [zhu2024fsgs] adapts 3DGS to sparse view synthesis through an efficient Gaussian expansion strategy, incorporating pretrained monocular depth estimation as a geometric prior to guide reconstruction. Similarly, DNGaussian [li2024dngaussian] exploits depth information by employing a hard and soft depth‑regularization scheme alongside a global‑local depth‑normalization strategy to reshape Gaussian primitives for more accurate geometry. Moreover, LoopSparseGS [bao2025loopsparsegs] proposes a loop-based framework that progressively densifies Gaussian initialization, aligns both scale-consistent and inconsistent depths for reliable geometry, and adopts sparse-aware sampling to mitigate oversized ellipsoid artifacts.
These studies address the challenge of sparse view reconstruction from different perspectives, proposing a variety of strategies to enhance reconstruction quality. In summary, their methods mainly rely on incorporating additional cues such as depth priors or introducing regularization mechanisms to alleviate overfitting. However, the fundamental limitation of sparse view reconstruction lies in the inherent lack of supervisory signals caused by extreme data sparsity, which restricts cross-view generalization and compromises geometric consistency. Consequently, while these methods offer partial improvements in visual quality, they do not fundamentally resolve the core issue. This persistent challenge drives our curriculum-based regularization framework, which aims to tackle the root cause of supervision deficiency rather than merely mitigating its symptoms.
Motivated by the above observations, we propose CuriGS, a curriculum-guided 3DGS framework designed for sparse view 3D reconstruction. The key idea of CuriGS is the introduction of student views—pseudo-views generated around real cameras (teacher) with controllable perturbation magnitudes. During training, a curriculum schedule progressively unlocks student groups with larger perturbations, allowing the model to gradually adapt from locally consistent to more diverse viewpoints. At each iteration, a subset of student views is randomly sampled from the currently active group and optimized through depth-correlation and co-regularization constraints, which enforce geometric consistency while mitigating reconstruction errors. Meanwhile, CuriGS maintains the best-performing student for each teacher and perturbation group, evaluated via a multi-signal metric combining structural similarity (SSIM), perceptual similarity (LPIPS), and a no-reference image-quality score [agnolucci2024qualityaware]. Students that exceed a predefined quality threshold are periodically promoted into the training set, effectively augmenting sparse supervision with reliable pseudo-views. This design effectively enriches the supervision signal under sparse-view conditions, improving both geometric fidelity and rendering realism while enhancing the model’s generalization to unseen viewpoints. The main contributions of this work are as follows:
  1. 1.
    Curriculum-guided sparse view expansion. We introduce the first curriculum-guided 3DGS framework that dynamically generates and promotes student views, expanding supervision directly from sparse inputs while mitigating overfitting and geometric inconsistency.
  2. 2.
    Unified pseudo-view learning mechanism. CuriGS provides a principled framework for generating, evaluating, and integrating pseudo-views into scene optimization, opening a new direction for virtual-view learning in reconstruction and synthesis tasks.
  3. 3.
    Superior performance and generalization. Extensive experiments across multiple benchmarks demonstrate that CuriGS consistently surpasses state-of-the-art baselines in rendering fidelity, perceptual quality, and geometric consistency.
2 Related Work
2.1 NeRF-based Sparse View Synthesis
While NeRF achieves photorealistic results under dense multi-view supervision, its optimization becomes severely ill-posed in sparse-view settings, often leading to geometry collapse, texture ambiguity, and overfitting to the observed views. To alleviate these issues, several methods [jain2021putting, somraj2023vip, deng2022depth, niemeyer2022regnerf, somraj2024simple, wang2023sparsenerf, yang2023freenerf] have been proposed to adapt NeRF-style representations for few-shot reconstruction.
SparseNeRF [wang2023sparsenerf] introduces weak geometric supervision by distilling depth-ordering priors from pretrained monocular depth estimators or coarse sensor depth. By injecting such constraints into the volumetric optimization, SparseNeRF reduces geometric ambiguity and mitigates depth collapse when only a handful of views are available. However, its performance heavily depends on the quality and domain alignment of the external depth priors, which may mislead optimization when the monocular estimator is biased or out-of-domain. FreeNeRF [yang2023freenerf] approaches the problem from a spectral and occlusion-regularization perspective, imposing frequency and near-camera density constraints on the learned radiance field to suppress unsupported high-frequency components and spurious near-field densities that commonly arise under sparse supervision. This family of regularizers is appealing for its simplicity and independence from explicit external priors, but it requires careful hyperparameter tuning and may attenuate genuine high-frequency surface details in richly textured scenes. PixelNeRF [yu2021pixelnerf] extends NeRF into a conditional formulation by leveraging learned multi-view feature fusion. Through large-scale multi-scene pretraining, it improves few-shot generalization and cross-scene adaptability, yet still struggles with limited fidelity and unstable geometry when applied to novel domains that diverge from the training distribution. Despite these advances, sparse-view NeRFs are fundamentally constrained by their implicit volumetric representation. Consequently, achieving stable geometry and consistent appearance reconstruction under extremely sparse inputs remains a significant challenge, motivating the development of alternative representations and training paradigms that can inherently cope with data scarcity.
2.2 3DGS-based Sparse View Synthesis
Compared with NeRF, 3DGS offers superior runtime efficiency, interpretable geometric primitives, and direct manipulability, making it particularly appealing for interactive visualization and efficient optimization. Motivated by these strengths, several studies [li2024dngaussian, zhu2024fsgs, zhang2024cor, zheng2025nexusgs, park2025dropgaussian, paliwal2024coherentgs, liu2026sv, zhang2026usgs] have explored its potential for sparse view synthesis tasks with most achieving notably better performance than NeRF-based methods.
CoR-GS [zhang2024cor] introduces an ensemble-based strategy in which multiple 3DGS instances are trained jointly, using mutual inconsistencies in both point-level geometry and rendered appearance to identify and suppress unreliable splats through co-pruning. This co-regularization effectively reduces reconstruction artifacts without relying on explicit ground-truth geometry. However, the method increases training complexity and computational overhead, and its effectiveness diminishes in highly symmetric or extremely sparse scenes where ensemble disagreement becomes ambiguous. NexusGS [zheng2025nexusgs] introduces an epipolar-depth–guided framework that integrates geometric priors to guide Gaussian densification, pruning, and blending. However, it relies on additional optical flow information and performs poorly when training views have low overlap, which limits its applicability. DropGaussian [park2025dropgaussian] proposes a lightweight structural regularization strategy by randomly deactivating subsets of Gaussians during training. This dropout-like mechanism redistributes gradients toward under-optimized splats, mitigating overfitting and improving generalization to unseen views. Unlike methods that rely on external priors or ensemble models, DropGaussian is simple and computationally efficient, though its stochastic nature may lead to unstable convergence or suboptimal results in extremely sparse or unstructured scenes.
Beyond the representative works discussed above, other 3DGS-based sparse-view methods [tang2025spars3r, xiong2025sparsegs, wan2025s2gaussian, bao2025loopsparsegs] have made progress by incorporating semantic guidance, Gaussian super-resolution, or confidence-weighted sparse enhancements to stabilize optimization. Nevertheless, these approaches fundamentally build upon stronger regularization or auxiliary supervision and do not address the key challenge of sparse view reconstruction, namely the intrinsic scarcity of training viewpoints, thus remaining methodologically constrained. In contrast, we directly tackle the fundamental limitation of sparse training supervision by expanding the available viewpoints through a curriculum-guided pseudo-view learning strategy.
3 Method
Figure 2: The overall architecture of the CuriGS framework is shown in (A). The pipeline consists of three key stages: (1) student view generation, where pseudo-camera poses are sampled around teacher views with multiple perturbation magnitudes, as detailed in (B); (2) curriculum scheduling, which gradually unlocks perturbation levels during training to progressively expand viewpoint diversity; and (3) student view evaluation and promotion, where each candidate is scored using perceptual (LPIPS), structural (SSIM), and no-reference quality metrics. Only the best student at each perturbation level that passes the evaluation criteria is promoted to the training set, as illustrated in (C). This curriculum-guided process enhances geometric consistency and rendering fidelity under sparse supervision.
In this section, we present CuriGS, a curriculum-guided, data-centric sparse-view 3DGS framework that progressively expands the effective training view distribution through the generation, evaluation, and selection of pseudo-views. The overall pipeline of the proposed method is illustrated in Fig. 2.
3.1 Preliminaries: Gaussian Splatting
3DGS represents a scene as a set of anisotropic Gaussian primitives, each parameterized by position $\mathbf{\mu}\in\mathbb{R}^{3}$, covariance $\Sigma\in\mathbb{R}^{3\times 3}$, opacity $\alpha\in[0,1]$, and spherical harmonic (SH) coefficients for view-dependent color. Formally, each Gaussian defines a density function in 3D space:
$G(\mathbf{x})=\exp\left(-\tfrac{1}{2}(\mathbf{x}-\mathbf{\mu})^{\top}\Sigma^{-1}(\mathbf{x}-\mathbf{\mu})\right)$ (1)
The rendering process follows a rasterization pipeline, in which each Gaussian is splatted into screen space, composited according to depth order, and accumulated using alpha blending. Specifically, the pixel color $\mathbf{C}(\mathbf{p})$ at screen location $\mathbf{p}$ is obtained by front-to-back alpha compositing:
$\mathbf{C}(\mathbf{p})=\sum_{i=1}^{N}T_{i}\,\alpha_{i}\,\mathbf{c}_{i}(\mathbf{p}),\quad T_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (2)
where $\alpha_{i}$ denotes the opacity of the $i$-th Gaussian, $\mathbf{c}_{i}(\mathbf{p})$ is its color contribution, and $T_{i}$ is the transmittance term accounting for accumulated transparency from closer Gaussians.
3.2 Curriculum: Student View Generation and Scheduling
Figure 3: Visualization of student view generation. Examples of pseudo-camera poses with different perturbation magnitudes around teacher view.
Student view generation.
The key idea of our framework is the introduction of student views with the goal of augmenting supervision by generating novel viewpoints that remain geometrically consistent with the original camera distribution, while progressively increasing diversity during training. Specifically, we generate student views by perturbing the extrinsics of each teacher camera within a controlled range. Given a teacher camera with rotation $\mathit{R}$ and translation $\mathit{T}$, we first calculate its optical center $\mathit{C}=-\mathit{R}^{\top}\mathit{T}$. Random angular perturbations in yaw and pitch are then applied, sampled from a zero-mean Gaussian distribution with a standard deviation $\sigma$ (in degrees), together with a mild radial perturbation $\sigma_{\mathit{r}}$ along the viewing direction, both of which collectively simulate small camera displacements.:
$R^{\prime}$ $=R\,R_{\Delta}(\sigma),\quad C^{\prime}=C(1+\epsilon_{r}),\quad\epsilon_{r}\!\sim\!\mathcal{N}(0,\sigma_{r}^{2}),$ (3)
$\hskip 93.89418ptT^{\prime}=-R^{\prime}C^{\prime}.$
The perturbed parameters $(R^{\prime},T^{\prime})$ define a new pseudo-view centered near the teacher pose and preserving identical intrinsic calibration.
For each teacher view $\mathcal{C}_{t}$, we initialize multiple groups of student views $\mathcal{P}(\sigma_{i})=\{\mathcal{C}_{s}^{\sigma_{i},j}\}_{j=1}^{N_{i}}$ with varying angular perturbation levels $\sigma_{i}$ drawn from a predefined range depending on the scene scale and sparsity. This hierarchical sampling naturally forms the basis of our curriculum strategy, where views with smaller perturbations are initially emphasized for stability, and those with larger perturbations are progressively unlocked as training proceeds. Fig. 3 illustrates examples of student views generated at different perturbation levels.
Curriculum scheduling.
Instead of introducing all student views simultaneously, we adopt a staged curriculum strategy. Training begins with students at small $\sigma$, ensuring stability by augmenting the dataset with near-teacher views that preserve local geometry. After a fixed number of iterations, the curriculum unlocks the next perturbation level, and this process continues iteratively until the largest $\sigma$ is reached, progressively exposing the model to increasingly diverse viewpoints.
We formalize the active perturbation level at iteration $t$ as:
$\sigma_{\text{active}}(t)=\min\big(\sigma_{\max},\,\sigma_{\min}+k\cdot\lfloor t/T_{s}\rfloor\big),$ (4)
where $\sigma_{\min}$ and $\sigma_{\max}$ denote the minimum and maximum angular perturbations, respectively. $T_{s}$ is the interval iterations after which the next perturbation level is unlocked, and $k$ is the increment step between successive stages.
By controlling the perturbation schedule through $\sigma_{\text{active}}(t)$, the model first consolidates local consistency around ground-truth views before gradually adapting to larger viewpoint variations. This design alleviates overfitting under sparse supervision and ensures a smooth training trajectory from local refinement to global generalization.
3.3 Student View Evaluation and Promotion
After generating student views, we design an evaluation-and-promotion mechanism that assesses their quality during training and selectively integrates the most reliable candidates into the training set. The purpose of this mechanism is to fully exploit the potential of student views, ensuring that only informative and geometrically consistent samples contribute to model learning.
Evaluation during training.
At each training iteration, for each teacher view $\mathcal{C}_{t}$, one student view $\mathcal{C}_{s}^{\sigma_{\text{active}},j}\in\mathcal{P}(\sigma_{\text{active}})$ is randomly sampled from the pool of student views corresponding to the currently active perturbation level $\sigma_{\text{active}}$. The corresponding rendered image is then compared against the teacher’s reference image through a composite multi-signal metric. This metric is designed to provide an objective evaluation of student views in the absence of ground-truth references, jointly accounting for structural similarity (SSIM) to assess spatial fidelity, perceptual distance (LPIPS) to measure feature-level realism, and a no-reference image-quality measure [agnolucci2024qualityaware] to capture global perceptual quality. For each $(\mathcal{C}_{t},\sigma_{i})$ pair, we maintain the best-performing student, defined as the candidate with the lowest evaluation loss up to the current iteration.
In addition to this scoring mechanism, student views are optimized using auxiliary pseudo-losses based on depth-correlation and co-regularization. Unlike the visual scoring, which serves to select among candidate views, these regularizers operate directly on the selected student views to enhance geometric alignment and multi-view consistency. Further details of these regularization terms are provided in Sec. 3.4.
Promotion to training views.
As the curriculum advances to a new perturbation level $\sigma_{\text{next}}$, the best-performing student view retained from the previous level $\sigma_{\text{prev}}$ for each teacher is subsequently evaluated. If its visual quality score exceeds a predefined threshold, the corresponding student view is then promoted to the official training set as a valid camera pose. In this manner, only high-quality, geometrically consistent pseudo-views are incorporated, gradually expanding the training coverage.
3.4 Losses and Training Objective
Our overall training objective consists of three complementary parts: (1) a photometric loss computed over all current training views, including the promoted student views; (2) a photometric loss on the original ground-truth views; and (3) the student-view regularization term.
Training and Ground-Truth Views.
While newly promoted student views effectively expand the sparse training distribution, their pseudo-supervision may inevitably introduce uncertainty. To prevent potential drift and maintain stable optimization, we additionally preserve explicit supervision on the original teacher views throughout training. This dual-level formulation allows the network to continuously refine its geometry and appearance while anchoring the learning process to reliable real observations. We employ a hybrid photometric loss that combines $\mathcal{L}_{\text{1}}$ and D-SSIM:
$\mathcal{L}_{\text{recon}}=(1-\lambda_{s})\mathcal{L}_{\text{1}}+\lambda_{s}\mathcal{L}_{\text{ssim}},$ (5)
Student Views.
Since student views are not real camera observations, their rendered outputs lack ground-truth images for direct supervision. Therefore, we introduce two pseudo-supervision schemes to optimize the learning of student views. The first scheme is based on depth-correlation. Specifically, we employ a pretrained monocular depth estimation model [ranftl2021vision] to predict a pseudo-depth map from the rendered image of a student view. This pseudo-depth is then compared with the depth map directly output by the 3DGS rasterization. To quantify their structural alignment, we adopt the Pearson correlation to formulate a relaxed relative loss. The formula:
$\mathcal{L}_{\text{depth}}^{s}=1-\frac{\mathrm{Cov}(D_{r},D_{p})}{\sigma(D_{r})\,\sigma(D_{p})},$ (6)
The second scheme relies on a dual-model consistency constraint [zhang2024cor]. During training, we maintain two 3DGS models with independent parameters, rendering the same student view in parallel. The photometric loss between their rendered results serves as a self-supervised constraint, enforcing cross-model visual consistency and reducing uncertainty introduced by view perturbations. We define the final student loss as:
$\mathcal{L}_{\text{stu}}=\lambda_{d}\,\mathcal{L}_{\text{depth}}^{s}+\lambda_{p}\,\mathcal{L}_{\text{photo}}^{s}.$ (7)
Overall Objective.
The overall optimization objective is formulated as:
$\mathcal{L}=\lambda_{\text{1}}\,\mathcal{L}_{\text{train}}+\lambda_{\text{2}}\,\mathcal{L}_{\text{gt}}+\lambda_{\text{3}}\,\mathcal{L}_{\text{stu}},$ (8)
4 Experiments
Table 1: Quantitative comparison with state-of-the-art methods on the LLFF, MipNeRF-360, and DTU datasets. CuriGS achieves the best or comparable results across PSNR, SSIM, and LPIPS metrics under sparse-view conditions.
Method LLFF (3 Views) MipNeRF-360 (24 Views) DTU (3 Views)
[…]
RegNeRF[niemeyer2022regnerf] 19.08 0.587 0.336 22.19 0.643 0.335 18.89 0.745 0.190
SparseNeRF[wang2023sparsenerf] 19.86 0.624 0.328 22.85 0.693 0.315 19.55 0.769 0.201
FreeNeRF[yang2023freenerf] 19.63 0.612 0.308 22.78 0.689 0.323 19.92 0.787 0.182
3DGS[kerbl20233d] 18.54 0.588 0.272 21.71 0.672 0.248 17.65 0.816 0.146
DNGaussian[li2024dngaussian] 19.12 0.591 0.294 18.06 0.423 0.584 18.91 0.790 0.176
FSGS[zhu2024fsgs] 20.43 0.682 0.248 23.40 0.733 0.238 17.14 0.818 0.162
CoR-GS[zhang2024cor] 20.45 0.712 0.196 23.55 0.727 0.226 19.21 0.853 0.119
DropGaussian[park2025dropgaussian] 20.76 0.713 0.200 23.66 0.747 0.233 19.31 0.857 0.142
NexusGS[zheng2025nexusgs] 21.00 0.730 0.179 23.86 0.753 0.206 20.21 0.869 0.102
LoopSparseGS[bao2025loopsparsegs] 20.85 0.717 0.205 24.09 0.755 0.226 20.68 0.856 0.125
Ours 21.10 0.732 0.193 24.10 0.759 0.205 20.35 0.870 0.131
Figure 4: Qualitative comparison on the LLFF dataset. CuriGS achieves sharper details and reduced texture drift compared with other baselines.
4.1 Datasets & Baselines
Datasets.
We evaluated our method on three widely adopted datasets: LLFF [mildenhall2019local], DTU [jensen2014large], and Mip-NeRF360 [barron2022mip]. Across all datasets, we assume that the camera parameters are known. All experimental settings, including data splits and image resolutions, are kept consistent with previous studies [zhu2024fsgs, zheng2025nexusgs, park2025dropgaussian] to ensure fair comparison.
Baselines.
We compared our approach with a representative set of state-of-the-art methods from two categories, including NeRF-based approaches [niemeyer2022regnerf, wang2023sparsenerf, yang2023freenerf] and recent 3DGS-based methods [li2024dngaussian, zhu2024fsgs, zhang2024cor, zheng2025nexusgs, park2025dropgaussian]. For quantitative evaluation, we reported the standard image space metrics PSNR, SSIM, and LPIPS computed on held-out test views.
4.2 Quantitative and Qualitative Results
LLFF.
On forward-facing real-world scenes from the LLFF, CuriGS achieves notable gains in both pixel and perceptual level metrics, as shown in Tab. 1. In particular, CuriGS attains the highest PSNR (21.10 dB) and SSIM (0.732) among all compared 3DGS variants, while achieving LPIPS performance competitive with the top baselines. These quantitative improvements correspond to a consistent reduction in cross-view texture drift and localized blurring in visual comparisons (Fig. 4). CuriGS recovers sharper high-frequency appearance around edges and small structures where other methods typically smear or misalign texture. The improvement indicates that selective promotion of high-quality pseudo-views effectively supplements supervision around original camera viewpoints, stabilizing both photometric and geometric reconstruction on challenging forward-facing scenes.
Figure 5: Qualitative comparison on the MipNeRF-360 dataset. CuriGS demonstrates improved perceptual fidelity and geometric stability on large-scale, unbounded scenes.
Mip-NeRF360.
For large-scale, unbounded scenes from the MipNeRF-360, CuriGS demonstrates robust perceptual quality and strong generalization. Quantitatively, it achieves the best results among the tested methods, with a PSNR of 24.10 dB, SSIM of 0.759, and LPIPS of 0.205 (Tab. 1). The corresponding visual comparisons are shown in Fig. 5. Compared to other baselines, CuriGS exhibits reduced unnatural color shifts and more accurately reconstructs fine details, such as grass and ground textures. These improvements stem from the curriculum strategy, which progressively exposes the model to larger perturbations, while the promotion mechanism filters out inconsistent pseudo-views, ultimately yielding more coherent long-range geometry and perceptually plausible renderings in large-scale scenes.
Figure 6: Qualitative comparison on the DTU dataset. CuriGS better preserves fine geometric structures and thin details compared to baseline methods under 3-views sparse supervision.
DTU.
On the object-centric DTU benchmark under a 3-views sparse setup, CuriGS shows clear advantages in geometry-sensitive metrics, achieving the highest SSIM, along with competitive PSNR and LPIPS, as shown in Tab. 1. Qualitative comparisons in Fig. 6 further highlight that CuriGS better preserves the overall structural integrity of objects while more effectively maintaining thin structures and fine geometric details, such as small protrusions and edges, which are often lost or distorted by other methods.
4.3 Ablation Study
To better quantify and validate the effectiveness of the proposed method in sparse view reconstruction, we conducted a comprehensive ablation study. Specifically, the experiments focused on two aspects: (1) the performance of the method when the curriculum guidance is removed, and (2) the influence of different numbers of initial training views on reconstruction performance. Since incorporating a depth loss on training views has been widely demonstrated to be beneficial and is not the primary contribution of this study, this component is retained in our baseline model. In the experimental setup, we selected representative samples with typical scene characteristics from each dataset for testing, rather than using all available data.
As reported in Tab. 2, the proposed curriculum guidance leads to consistent improvements in PSNR, SSIM, and LPIPS across all datasets and view configurations.
On the LLFF dataset, curriculum guidance consistently improves performance under all sparse-view conditions. When trained with only two input views, the full strategy achieved a PSNR of 20.94 dB compared to 18.30 dB without guidance, alongside notable increases in SSIM (0.768 vs. 0.601) and decreases in LPIPS (0.194 vs. 0.231). As the number of views increased, overall performance improved, but the model without curriculum guidance still lagged behind.
A similar pattern was observed on the MipNeRF-360 dataset. With 16 views, the inclusion of curriculum guidance improved PSNR from 20.32 dB to 21.23 dB and SSIM from 0.754 to 0.781, while LPIPS decreased from 0.202 to 0.189. At 24 views, the improvements remained consistent, indicating that the proposed mechanism maintains strong generalization capability even as the complexity of the scene increases.
The improvements were even more pronounced on the DTU dataset, which contains structured indoor objects with rich geometric details. The full strategy yielded 3.5–4 dB higher PSNR compared with the baseline lacking curriculum guidance. For example, at three training views, PSNR increased from 18.46 to 22.65, SSIM improved from 0.922 to 0.947, and LPIPS dropped from 0.065 to 0.050.
Table 2: Ablation study on the effect of curriculum guidance under different view numbers across the LLFF, MipNeRF-360, and DTU datasets.
Dataset #Views / Setting PSNR SSIM LPIPS
LLFF 2 / Full 20.94 0.768 0.194
2 / w/o Cur. 18.30 0.601 0.231
3 / Full 22.51 0.834 0.150
3 / w/o Cur. 21.03 0.772 0.202
5 / Full 23.95 0.861 0.140
5 / w/o Cur. 22.97 0.841 0.143
MipNeRF-360 16 / Full 21.23 0.781 0.189
16 / w/o Cur. 20.32 0.754 0.202
24 / Full 23.43 0.851 0.131
24 / w/o Cur. 22.91 0.830 0.142
DTU 2 / Full 18.66 0.922 0.077
2 / w/o Cur. 15.04 0.896 0.092
3 / Full 22.65 0.947 0.050
3 / w/o Cur. 18.46 0.922 0.065
6 / Full 24.45 0.947 0.043
6 / w/o Cur. 20.51 0.946 0.048
5 Conclusion
In this work, we present CuriGS, a curriculum-guided framework for sparse view 3D reconstruction using 3DGS. Our method tackles the core challenge of sparse supervision by introducing a hierarchical student-view learning mechanism, in which pseudo-views are progressively sampled and regularized according to a curriculum schedule. By leveraging multi-signal evaluation metrics that combine perceptual, structural, and image-quality measures, CuriGS selectively promotes high-quality student views, thereby enhancing geometric consistency and photometric fidelity. Extensive experiments on the LLFF, MipNeRF-360, and DTU datasets demonstrate that CuriGS consistently outperforms existing NeRF- and 3DGS-based baselines under various sparse-view settings. The curriculum-guided strategy substantially mitigates overfitting, improves generalization to unseen viewpoints, and preserves fine geometric details even under extreme sparsity.
[…]
Supplementary Material
6 Implementation Details
6.1 Dataset Setup
LLFF and MipNeRF-360.
Following standard protocols, we use the LLFF and MipNeRF-360 datasets for evaluating forward-facing and unbounded real-world scenes, respectively. For each dataset, every 8th image is selected as part of the test set, and the remaining images are used for training. To simulate sparse-view conditions, we uniformly subsample a limited number of training views from the available set. All images are downsampled by a factor of 8 for computational efficiency and consistency with previous baselines [zhu2024fsgs, zheng2025nexusgs, park2025dropgaussian, li2024dngaussian, zhang2024cor].
DTU.
Following prior work, we select 15 representative scenes (IDs 8, 21, 30, 31, 34, 38, 40, 41, 45, 55, 63, 82, 103, 110, and 114) from the DTU dataset. For each scene, views 25, 22, and 28 are used as the 3-view training set, while the remaining 25 views are designated for testing. All images are downsampled by a factor of 4, and binary masks are applied to remove background regions, ensuring that both training and evaluation focus solely on the target objects.
6.2 Training Details
LLFF.
For the LLFF scenes, the model is trained for 30,000 iterations. The curriculum-based student sampling phase starts at iteration 3,000 and ends at iteration 24,000. For each teacher view, we generate pseudo-student views using perturbation levels $\sigma\in\{1,2,3,4,5,6,7,8,9,10\}$, with 5 students per level. The quality threshold is set to 0.4.
MipNeRF-360.
In the MipNeRF-360, for the bicycle and stump scenes, the curriculum phase spans from iteration 10,000 to 30,000, with a total of 33,000 iterations. For the remaining scenes, training runs for 30,000 iterations, with the curriculum activation scheduled between iterations 7,000 and 27,000. Student views are generated using perturbation levels $\sigma\in\{2,4,6,8,10\}$, with 3 students per level. The quality threshold is set to 0.45.
DTU.
For the DTU dataset, the model is trained for 13,000 iterations, and curriculum occurs between iterations 2,000 and 12,000. We generate student views using perturbation levels $\sigma\in\{1,2,3,4,5\}$, with 10 students per level. The quality threshold is set to 0.55 for scans 114, 110, and 103, and 0.45 for the remaining scans.
Similar to teacher views, student views are also processed with background masking to exclude irrelevant regions. To obtain accurate masks for pseudo views, we first identify background pixels in each corresponding teacher image. Let $\mathcal{B}_{t}$ denotes the set of background pixel coordinates in the teacher image, and let $\mathbf{I}_{s}(p)$ represents the RGB color of pixel $p$ in the student view. For pixels at corresponding positions in the student view, we compute the color mean $\boldsymbol{\mu}_{s}$ and variance $\boldsymbol{\sigma}_{s}^{2}$ over $\mathcal{B}_{t}$. A student pixel $p$ is classified as background if its color lies within a fixed deviation range around the teacher’s background statistics:
$M_{s}(p)=\begin{cases}1,&\text{if }\left\lVert\mathbf{I}_{s}(p)-\boldsymbol{\mu}_{s}\right\rVert<\tau\,\boldsymbol{\sigma}_{s},\\ 0,&\text{otherwise},\end{cases}$ (9)
where $M_{s}(p)$ denotes the binary background mask for the student view, and $\tau$ is a constant threshold controlling the tolerance range. This adaptive masking strategy enables robust background suppression and maintains mask consistency despite geometric perturbations between teacher and student cameras.
7 Extended Ablation Analysis
Figure 7: Evolution of PSNR, SSIM, and LPIPS during training on a DTU scene with and without curriculum guidance.
We previously demonstrated that the introduction of student views effectively alleviates overfitting in sparse-view reconstruction. To further quantify this effect, we conducted a detailed training analysis on a representative DTU scene, continuously monitoring the evolution of PSNR, SSIM, and LPIPS throughout the optimization process for both the full CuriGS framework and an ablated version without curriculum guidance, as shown in Fig. 7. As training progresses, the model without curriculum guidance exhibits clear signs of overfitting—its performance metrics gradually deteriorate as it memorizes limited viewpoints rather than improving scene generalization. In contrast, the full CuriGS strategy maintains steady performance gains across all metrics, particularly in the later training stages, demonstrating that the curriculum-based student-view integration provides sustained and stable supervision that effectively prevents overfitting and promotes consistent convergence.
8 Discussion
Figure 8: Visualization of a teacher view and its associated student-view group generated with an example perturbation magnitude. From left to right, the rendered results of the students are shown at different training stages, demonstrating how visual fidelity and structural coherence improve over time. Scores are computed as a combination of multiple metric losses, with lower values indicating better quality The best student is labeled with a blue box.
The motivation behind our curriculum-based learning strategy is to effectively expand the sparse set of training views. In sparse-view scenarios, limited camera coverage restricts supervision diversity, making it crucial to introduce additional, yet reliable, viewpoints. Therefore, a key challenge lies in how to efficiently select and utilize high-quality student views. To this end, we design a progressive training mechanism where student views with different perturbation magnitudes are gradually introduced to assist optimization, and the best-performing students are continuously tracked for potential inclusion in the training set. To evaluate student quality, we integrate three complementary metrics—LPIPS, SSIM, and a no-reference image-quality score—forming a unified criterion that compensates for the absence of explicit ground-truth supervision in pseudo views. LPIPS and SSIM capture perceptual and structural similarity with the corresponding teacher render, ensuring semantic and visual consistency, while the no-reference score enforces overall image plausibility. Empirically, this composite evaluation yields a stable and balanced metric for identifying reliable student views, effectively bridging the supervision gap between real and synthetic viewpoints.
When training reaches a perturbation-switching stage, the best student from each group undergoes a final quality evaluation. If its rendered result surpasses a predefined threshold, the view is promoted into the training set. At this stage, only the image-quality score is used for evaluation, since LPIPS and SSIM tend to favor students closer to the teacher view and thus provide limited diversity. Allowing moderate geometric or appearance deviations encourages the model to explore broader scene variations, ultimately improving generalization and robustness. Fig. 8 visualizes a teacher view and its corresponding student group at a specific perturbation level, showing progressive refinement of student renders and the selection of the best-performing candidate.
9 Additional Visualization Results
To further illustrate the effectiveness and robustness of our approach, we provide additional qualitative comparisons on the LLFF, Mip-NeRF360, and DTU datasets. As shown in Figs. 9, 10 and 11, CuriGS produces visually coherent and geometrically stable reconstructions across diverse scene types, ranging from real forward-facing captures to unbounded outdoor environments and object-centric setups. These results collectively demonstrate the versatility and generalization ability of CuriGS across different domains and levels of sparsity.
Figure 9: Per-scene visualization results on LLFF.
Figure 10: Per-scene visualization results on MipNeRF-360.
Figure 11: Per-scene visualization results on DTU.